OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 2x + y = 3 -3x + 4y = 1 I forgot a lot about math over summer, can someone please explain to me how to do this?

OpenStudy (anonymous):

first from the equation 3x + 4y = 1, solve for the value of x..can u do it..?

OpenStudy (anonymous):

-3x + 4y = 1 rather...

OpenStudy (anonymous):

Would I go -3x = 1 - 4y?

OpenStudy (anonymous):

Alternatively solve for y from the firts equation that may be comapartively easier.. from 1st eq. we have y = 3-2x

OpenStudy (anonymous):

yes matricked is right,.. i thing its more easy...

OpenStudy (anonymous):

oh ok i see, and then what would I do from there? I am not the best at this sorta thing.

OpenStudy (anonymous):

now plug in 3-2x in place of y in the 2nd eq. -3x+4y=1 after substituting we have -3x +4 (3-2x) = 1 ihope you can do the rest now...

OpenStudy (anonymous):

a. (1, 1) b. (-1, -1) c. (1, -1) d. (-1, 1) the multiple choice says this, if can someone please explain to me what the answer is and why that would be very helpful.

OpenStudy (anonymous):

-3x +4 (3-2x) = 1 or -3x +12 -8x = 1 or -3x -8x = 1-12 -11x = -11 henc x = 1 now we need y we can obtain it from y=3-2x as we know x=1 so y=3 -2(1)=3-2=1 thus the solution is x=1 and y=1 in the choice the first number is for x and the last one for y so here we have a. (1, 1)